The length of the clocks second hand is 45 mm. What's the average velocity of the end point of this second hand during any thirty seconds of time?
1
Expert's answer
2019-01-16T11:47:11-0500

The average velocity of the end point of the second hand

v=s/t

The displacement of the end point of the second hand during any thirty seconds

s=2R

So

v=2R/t

=(2×45 mm)/(30 s)=3 mm/s